95k:20084 20J06

Adem, Alejandro ^{(}1-WI^{)}; Milgram, R. James ^{(}1-STF-HT^{)} Invariants and cohomology of groups. Papers in honor of José Adem (Spanish). Bol. Soc. Mat. Mexicana (2) 37 (1992), no. 1-2, 1–25. Calculation of the mod p cohomology of a finite group often depends on the calculation of rings of invariants for the action of normalizers on the cohomology of elementary abelian p-subgroups. This paper contains a general discussion along these lines, and some explicit calculations of invariant rings for various subgroups of GL(4, F_{2}) ≅ A_{8} acting on a rank-four elementary abelian 2-group. These calculations play a role in the calculation of the mod 2 cohomology of the sporadic simple groups J_{2}, J_{3}, M_{22}, M_{23} and O’N. David Benson (1-GA)

\begin{matrix}c \\ ◯ \end{matrix} Copyright American Mathematical Society 1995, 1996